# Enhancing Image Layout Control with Loss-Guided Diffusion Models

### v1

@@ 4.2 Loss Guidance
This no longer corresponds to sampling from an approximation to $\hat{q}(\mathbf{z}_{0}|\mathbf{y})$ (see Song et al. (2023)), however, so long as the latents $\mathbf{z}_{t}$ are not too out-of-distribution with respect to the marginals $q(\mathbf{z}_{t}|\mathbf{y})$, then, much like CFG Ho and Salimans (2021), this process influences the trajectory to favor samples from $p_{\theta}(\mathbf{z}_{0}|\mathbf{y})$ for which the loss term is small.
@@
$\ell_{\mathbf{y}}(\mathbf{z}_{t})=\sum_{j\in S}g_{\text{IB}}(\mathbf{\bar{m}}_{j}\odot(A_{t})_{j})-g_{\text{OB}}(\mathbf{m}_{j}\odot(A_{t})_{j}),$ (15)
where $\mathbf{m}_{j}$ is a mask whose value is $1$ over the region where token $\mathbf{y}_{j}$ should appear, and 00 otherwise, and $\mathbf{\bar{m}}_{j}=1-\mathbf{m}_{j}$. Intuitively, this simple loss encourages sampling latents whose attention maps for each target token take their largest values within the masked regions.
The specific choices of $g_{\text{IB}}$ and $g_{\text{OB}}$ heavily influence the behaviour of loss-guidance. In BoxDiff, Xie et al. (2023) use sums over the $P\ll n$ most attended-to pixels in the masked attention maps. This seems to be an essential step for maintaining good image quality at very high levels of loss guidance. Since we will use much lower levels of loss guidance, we find that choosing $g_{\text{IB}}=g_{\text{OB}}=\text{sum}(\cdot)$ works well to contain the attentions within the regions defined by $m$.
In practice, diffusion models are trained to predict the normalized noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ in $\mathbf{z}_{t}$. Revisiting Eq. (7), we observe that we can define the corresponding modified noise prediction by scaling the loss-guidance term appropriately Dhariwal and Nichol (2021):
@@
Figure 3: A graphical depiction of injection loss guidance (iLGD).
@@
A weakness of loss-guidance is that the ad-hoc choice of the loss function may not compete well with the predicted score $\mathbf{s}_{\theta}(\mathbf{z}_{t},t)$. In each denoising step, the latent $\mathbf{z}_{t-1}$ must fall in a high probability region of of $q(\mathbf{z}_{t-1}|\mathbf{z}_{t})$, otherwise the predicted latent $\mathbf{z}_{t-1}$ will be out-of-distribution, which results in degraded image quality in the final sampled latent $\mathbf{z}_{0}$. This means that the guidance term’s influence in Eq. (12) should be small enough to avoid moving the latents into low probability regions. On the other hand, small loss guidance strengths may exert too little influence on the sampling trajectory. In this case, the model produces in-distribution samples, but ones which do not fully agree with the desired layout. This tradeoff is illustrated in the first row of Figure 1. One advantage of loss-guidance is that, even at low strengths, it is able to exert some influence over sampling trajectories without biasing the style of the images.
@@
Construct the mask $m=\{\mathbf{m}_{1},\mathbf{m}_{2},\ldots,\mathbf{m}_{k}\}$ from the bounding boxes;
@@
      $\ell_{\mathbf{y}}(\mathbf{z}_{t})=\begin{cases}\sum_{j\in S}g_{\text{IB}}(\mathbf{\bar{m}}_{j}\odot(A_{t})_{j})-g_{\text{OB}}(\mathbf{m}_{j}\odot(A_{t})_{j})&\text{if $t>t_{\text{loss}}$},\\ 0&\text{otherwise}\end{cases}$;
@@ 5.1 Experimental Setup
We follow Xie et al. (2023) and evaluate performance on a dataset consisting of 200 prompt and bounding box pairs, spanning 20 different prompts and 27 object categories. Each prompt reflects either of the following prompt structures: “a {} …,” or “a {} and a {} ….”
@@
To evaluate each method’s faithfulness to the prescribed bounding box, we use YOLOv4 Bochkovskiy et al. (2020) to compare the predicted bounding boxes over the set of generated images to the ground truth bounding boxes, and report the average precision at $\text{IOU}=0.5$.
@@
More details about these evaluation metrics can be found in Appendix C.
5.2 Implementation Details
We implemented our method on the official Stable Diffusion v1.4 model  Rombach et al. (2022) from HuggingFace. All images are generated using 50 denoising steps and a classifier-free guidance scale of 7.5, unless otherwise noted. We use the noise scheduler LMSDiscreteScheduler Karras et al. (2022) provided by HuggingFace. Experiments were conducted on an NVIDIA TESLA V100 GPU.
When attempting to control the layout of a generated image, we find that skipping the first step, so that it remains a standard denoising step, leads to better results. We do this for all experiments conducted in this paper which use either injection or loss guidance or both. We use $\eta=0.48$, $\nu^{\prime}=0.75$, $t_{\text{loss}}=25$, and $t_{\text{inject}}=10$, unless otherwise noted.
In our comparisons with BoxDiff, we maintain the default parameters the authors provide in their implementation. We start with $\alpha_{T}=20$, which decays linearly to $\alpha_{0}=10$, and perform guidance for 25 denoising steps out of a total of 50 steps.
@@ 5.3 Comparisons
Figure 4: A comparison of iLGD against BoxDiff and Stable Diffusion. The random seed is kept the same across each set of images.
Table 1: Comparison of the average contrast and saturation over 200 images for BoxDiff, Stable Diffusion (SD), and iLGD.
@@
SD + Injection 46.07 105.63
SD + Loss Guidance 58.51 102.62
Table 2: Comparison of various quality metrics for BoxDiff, Stable Diffusion (SD), and iLGD, averaged over 200 images.
@@
We present a comparison between our proposed method (iLGD), BoxDiff  Xie et al. (2023), and Stable Diffusion Rombach et al. (2022) in Figure 4. Qualitatively, we observe that BoxDiff produces images with much higher contrast and saturation. For instance, for the prompt “a donut and a carrot,” both objects in BoxDiff’s image appear unnaturally bright or dark in certain regions. Similarly, Figure 1 indicates that BoxDiff produces images with very high contrast, e.g., the light shade of the bread and the dark shade of the bottle, and saturation, e.g., the red of the tomatoes, when using high loss guidance strengths to produce the desired layout. High contrast and saturation are even visible in some images when just Stable Diffusion is used, e.g., in the prompts “a balloon and a cake and a frame…” and “a suitcase and a handbag” in Figure 5.
These observations are reflected quantitatively in Table 1, which reports the average contrast and saturation across all generated images. The same high-contrast and high-saturation phenomenon observed in BoxDiff is seen to occur in Stable Diffusion when increasing the classifier-free guidance (CFG) scale. Both loss guidance and CFG appear to move the predicted latent into similar low-probability regions if the strength is too large, which may nonetheless be necessary to obtain either the appropriate layout or the appropriate agreement with the semantics of the text prompt. Attention injection does not appear to have this biasing effect, and the contrast and saturation for iLGD are similar to those of attention injection.
@@
Biasing the latents using injection also clearly helps to preserve image quality and achieve stronger layout control, particularly for layouts which are difficult for the model to generate. For the prompt “a red book and a clock,” BoxDiff struggles to move the clock into the correct position, instead generating it as a lower quality, shadow-like figure. Using iLGD, the clock maintains its fine details, e.g., the numbering on the face, as well as its texture. For the prompt “a bowl and a spoon,”, the image produced by iLGD is more faithful to the bounding boxes, whereas BoxDiff has difficulty reorienting the scene. The bowl remains in the same position as it was with Stable Diffusion alone, and it becomes difficult to distinguish the spoon from its shadow as dark colours are exaggerated.
In Table 2, we report various metrics related to image quality (CLIP-IQA), bounding box precision (YOLOv4), and text-to-image similarity (T2I-Sim). While both BoxDiff and iLGD achieve similar T2I-Sim and AP@0.5 scores, the latter achieves better performance on all CLIP-IQA metrics. This indicates that iLGD is able to achieve superior image quality without sacrificing layout control, qualifying it as a meaningful improvement over BoxDiff.
We provide additional comparisons in Appendix B.
@@ 5.4 Ablation Studies
To understand the effectiveness of iLGD, we compare it against injection alone. Figure 5 presents a visual comparison between the two methods. We observe that, while injection typically does generate an object in each bounding box, the object itself may be incorrect. To illustrate, for the prompt “a balloon, a cake, and a frame, …,” using injection leads to a frame appearing where the cake should be; for “a donut and a carrot,” it generates a hand where the donut should be; for “a suitcase and a handbag,” a second suitcase is generated instead of a handbag; and for “a cat sitting outside,” something akin to a tree stump appears instead of a cat. In the example for “a castle in the middle of a marsh,” the castle does not fill up the bounding box, and for “a cat with a tie” the resulting image has the correct layout, but begins to appear like a cartoon cutout. When loss guidance is added using iLGD, all of these images appear correctly. These observations are reflected in Table 2, where iLGD achieves a noticeably higher AP@50 score compared to injection.
@@ A.1 Denoising Diffusion Probabilistic Models
In denoising diffusion probabilistic models (DDPM) Ho et al. (2020), the forward process is characterized by the Markov chain $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})\sim\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$, for some noise schedule $\beta_{t}$. In this case, $q(\mathbf{x}_{t}|\mathbf{x}_{0})\sim\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ and $\alpha_{t}=1-\beta_{t}$. The reverse process is typically modeled by a learned Markov chain $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\sim\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\sigma_{t}^{2}\mathbf{I})$, where $\sigma_{t}^{2}=\beta_{t}$.
@@
The right hand side is the usual evidence lower bound (ELBO), which is minimized instead. In Ho et al. (2020), the authors show that minimizing the ELBO is equivalent to minimizing
@@
for some $\lambda(t)$, where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $\mathbf{x}_{t}(\mathbf{x}_{0},\epsilon)=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, and

### v2

@@ 4.2 Loss Guidance
This no longer corresponds to sampling from an approximation to $\hat{q}(\mathbf{z}_{0}|\mathbf{y})$ (see Song et al. (2023)), however, so long as the latents $\mathbf{z}_{t}$ are not too out-of-distribution with respect to the marginals $q(\mathbf{z}_{t}|\mathbf{y})$, then this process influences the trajectory to favor samples from $p_{\theta}(\mathbf{z}_{0}|\mathbf{y})$ for which the loss term is small.
@@
$\ell_{\mathbf{y}}(\mathbf{z}_{t})=\sum_{j\in S}\text{sum}(\mathbf{\bar{m}}_{j}\odot(A_{t})_{j})-\text{sum}(\mathbf{m}_{j}\odot(A_{t})_{j}),$ (9)
where $\mathbf{m}_{j}$ is a mask whose value is $1$ over the region where token $\mathbf{y}_{j}$ should appear, and 00 otherwise, and $\mathbf{\bar{m}}_{j}=1-\mathbf{m}_{j}$. Intuitively, this simple loss encourages the sampling of latents whose attention maps for each target token take their largest values within the masked regions.
When high levels of loss-guidance are used, the specific choice of the loss function heavily influences the behavior of the denoising process. In BoxDiff, Xie et al. Xie et al. (2023) use sums over the $P\ll n$ most attended-to pixels in the masked attention maps, while in Chen et al. Chen et al. (2024) the authors use a normalized loss which depends on the ratio $\text{sum}(\mathbf{m}_{j}\odot(A_{t})_{j})/\text{sum}((A_{t})_{j})$. These choices seem to be essential for enabling those methods to maintain good image quality at very high levels of loss guidance. Since we will use much lower levels of loss guidance, we find that our method is significantly less sensitive to the choice of loss function, with the result that our simple loss function works well to contain the attentions within the regions defined by $\mathbf{m}$.
In practice, diffusion models are trained to predict the scaled noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ in $\mathbf{z}_{t}$. Revisiting Eq. (2), we observe that we can define the corresponding modified noise prediction by scaling the loss-guidance term appropriately Dhariwal and Nichol (2021):
@@
Since the loss function $\ell_{\mathbf{y}}(\mathbf{z}_{t})$ is a scalar function, computing its gradient with respect to $\mathbf{z}_{t}$ by reverse accumulation requires only a single sweep through the computational graph. As a result, the runtime and memory requirements of sampling using loss guidance are approximately twice those of sampling without loss guidance. This can be seen in, for example, Table 1 of Chen et al. Chen et al. (2024).
@@
A weakness of loss-guidance is that the ad-hoc choice of the loss function may not compete well with the predicted score $\mathbf{s}_{\theta}(\mathbf{z}_{t},t)$. In each denoising step, the latent $\mathbf{z}_{t-1}$ must fall in a high probability region of of $q(\mathbf{z}_{t-1}|\mathbf{z}_{t})$, otherwise the predicted latent $\mathbf{z}_{t-1}$ will be out-of-distribution, which results in degraded image quality in the final sampled latent $\mathbf{z}_{0}$. This means that the guidance term’s influence in Eq. (6) should be small enough to avoid moving the latents into low probability regions. On the other hand, small loss guidance strengths may exert too little influence on the sampling trajectory. In this case, the model produces in-distribution samples, but ones which do not fully agree with the desired layout. This tradeoff is illustrated in the first row of Figure 1, which shows images with unnaturally high contrast and saturation appearing at high loss-guidance strengths . One advantage of loss-guidance is that, even at low strengths, it is able to exert some influence over sampling trajectories without biasing the style of the images.
@@
Figure 3: A graphical depiction of injection loss guidance (iLGD).
@@
Construct the mask $\mathbf{m}=\{\mathbf{m}_{1},\mathbf{m}_{2},\ldots,\mathbf{m}_{k}\}$ from the bounding boxes;
@@
      $\ell_{\mathbf{y}}(\mathbf{z}_{t})=\begin{cases}\sum_{j\in S}\text{sum}(\mathbf{\bar{m}}_{j}\odot(A_{t})_{j})-\text{sum}(\mathbf{m}_{j}\odot(A_{t})_{j})&\text{if $t>t_{\text{loss}}$},\\ 0&\text{otherwise}\end{cases}$;
@@ 5.1 Experimental Setup
We follow Xie et al. Xie et al. (2023) and evaluate performance on a dataset consisting of 200 prompt and bounding-box pairs, spanning 20 different prompts and 27 object categories. Each prompt reflects either of the following prompt structures: “a {} …,” or “a {} and a {} ….” MultiDiffusion requires separate prompts for the foreground elements and for the background, which we manually created from our original prompts. For example, the prompt “a red book and a clock” became the two foreground prompts “a red book” and “a clock,” with a null background prompt, while “a giraffe on a field” became the foreground prompt “a giraffe” in the foreground and the background prompt “a field.”
Methods
In this section, we compare our proposed method (iLGD), BoxDiff Xie et al. (2023), Chen et al. Chen et al. (2024), MultiDiffusion Bar-Tal et al. (2023), and Stable Diffusion Rombach et al. (2022). We also include the results of an ablation study of our method, in which we use just injection (SD + I) or just loss guidance (SD + LG) alone. In order to allow loss guidance to exert sufficient influence over the final image layouts when it is used in isolation, we increase the loss-guidance strength above the level used in iLGD (see Appendix B for more details). All of the methods compared in this section used the official Stable Diffusion v1.4 model Rombach et al. (2022) from HuggingFace.
@@ Methods
To evaluate each method’s faithfulness to the prescribed bounding box, we use YOLOv4 Bochkovskiy et al. (2020) to compare the predicted bounding box over the set of generated images to the ground truth bounding boxes, and report the average precision at $\text{IOU}=0.5$.
@@
More details about these evaluation metrics can be found in Appendix B.
@@ 5.2 Comparisons
Figure 4: A comparison of iLGD against BoxDiff, Chen et al., MultiDiffusion, and Stable Diffusion. The random seed is kept the same across each set of images.
Table 1: Comparison of the average contrast and saturation over 200 images for BoxDiff, Chen et al., MultiDiffusion, Stable Diffusion (SD), with injection (SD + I) and with loss guidance (SD + LG), and iLGD.
@@
Chen et al. 62.12 98.56
MultiDiffusion 48.03 75.01
SD + I 46.07 105.63
SD + LG 58.51 102.62
Ours (iLGD) 47.53 105.81
Table 2: Comparison of various quality metrics for BoxDiff, Chen et al., MultiDiffusion, Stable Diffusion (SD), with injection (SD + I) and with loss guidance (SD + LG), and iLGD, averaged over 200 images.
@@
We present a comparison between our proposed method (iLGD), BoxDiff Xie et al. (2023), Chen et al. Chen et al. (2024), MultiDiffusion Bar-Tal et al. (2023), and Stable Diffusion Rombach et al. (2022) in Figure 4. Qualitatively, we observe that BoxDiff and Chen et al. both produce images with higher contrast, and BoxDiff produces especially saturated images. For instance, for the prompt “a donut and a carrot,” both objects in BoxDiff’s image appear unnaturally bright or dark in certain regions. Chen et al. produces high contrast images for the prompts “a dog on a bench” and “a banana and broccoli,” though with visibly less saturation than BoxDiff. High contrast and saturation are even visible in some images when just Stable Diffusion is used, e.g., in the prompts “a balloon and a cake and a frame…” and “a suitcase and a handbag” in Figure 5.
These observations are reflected quantitatively in Table 1, which reports the average contrast and saturation across all generated images. The same high-contrast phenomenon observed in BoxDiff and Chen et al. is seen to occur in Stable Diffusion when increasing the classifier-free guidance (CFG) scale. Furthermore, high staturation can also be seen in both BoxDiff and Stable Diffusion at high CFG scales. Both loss guidance and CFG appear to move the predicted latent into similar low-probability regions if the strength is too large, which may nonetheless be necessary to obtain either the appropriate layout or the appropriate agreement with the semantics of the text prompt. Neither attention injection nor MultiDiffusion appear to have this biasing effect, and the contrast and saturation for iLGD are similar to those of attention injection.
@@
Biasing the latents using injection also clearly helps to preserve image quality and achieve stronger layout control, particularly for layouts which are difficult for the model to generate. For the prompt “a red book and a clock,” BoxDiff struggles to move the clock into the correct position, instead generating it as a lower quality, shadow-like figure, while Chen et al. omits the clock entirely, and MultiDiffusion generates a chaotic scene with many artifacts. Using iLGD, the clock and book appear in the correct positions, and the clock maintains its fine details, e.g., the numbering on the face, as well as its texture. For the prompt “a bowl and a spoon,” the image produced by iLGD is more faithful to the bounding boxes than image produced by BoxDiff, in which it is also difficult to distinguish the spoon from its shadow as dark colours are exaggerated, as well as the image of Chen et al., which struggles to reposition the objects. The layout produced by MultiDiffusion is more accurate, but the image itself is not sensible. In general, it appears that BoxDiff performs very well, but produces images of unnaturally high contrast and saturation. The method of Chen et al. produces images of reasonable quality, but struggles to move the objects to the desired bounding boxes. MultiDiffusion appears to consistently make objects appear in the bounding boxes, but often produces chaotic scenes or scenes with cut-and-paste artifacts, in which the foreground images seem glued to an incompatible background (see the prompt “a banana and broccoli” for one such example).
In Table 2, we report various metrics related to image quality (CLIP-IQA), bounding box precision (YOLOv4) and text-to-image similarity (T2I-Sim). While both BoxDiff and iLGD achieve similar T2I-Sim and AP@0.5 scores, the latter achieves better performance on all CLIP-IQA metrics. While the method of Chen et al. achieves T2I-Sim and CLIP-IQA scores that are almost as good as those of iLGD, it achieves a much lower AP@0.5 score. MultiDiffusion has the highest AP@0.5 score of all methods considered, but it also has the lowest T2I-Sim and CLIP-IQA metrics (all except for Clear), showing that MultiDiffusion sacrifices an unacceptable amount of image quality in exchange for layout control. The ablation study (SD + I) using just injection shows good image quality, but substantially worse layout control than iLGD, while the ablation study (SD + L) using just loss guidance shows slightly worse image quality and much worse layout control than iLGD. These results indicate that iLGD is able to achieve superior image quality without sacrificing layout control, qualifying it as a meaningful improvement over BoxDiff, Chen et al., and MultiDiffusion.
We provide additional comparisons in Appendix C.
@@ 5.3 Ablation Studies
A visual comparison between iLGD and injection alone is presented in Figure 5. We observe that, while injection typically does generate an object in each bounding box, the object itself may be incorrect. To illustrate, for the prompt “a balloon, a cake, and a frame, …,” using injection leads to a frame appearing where the cake should be; for “a donut and a carrot,” it generates a hand where the donut should be; for “a suitcase and a handbag,” a second suitcase is generated instead of a handbag; and for “a cat sitting outside,” something akin to a tree stump appears instead of a cat. In the example for “a castle in the middle of a marsh,” the castle does not fill up the bounding box, and for “a cat with a tie” the resulting image has the correct layout, but begins to appear like a cartoon cutout. When loss guidance is added using iLGD, all of these images appear correctly. These observations are reflected in Table 2, where injection achieves a noticeably lower AP@0.5 score compared to iLGD.
@@ A.1 Denoising Diffusion Probabilistic Models
Diffusion models Ho et al. (2020) are characterized by two principle algorithms. The first is the forward process, wherein the data $\mathbf{x}_{0}$ is gradually corrupted by Gaussian noise until it becomes pure noise, which we denote by $\mathbf{x}_{T}$. The reverse process moves in the opposite direction, attempting to recover the data by iteratively removing noise. The denoiser $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is typically a UNet Ronneberger et al. (2015) which accepts an image $\mathbf{x}_{t}$, and predicts its noise content $\bm{\epsilon}$. Removing a fraction of this noise yields a slightly denoised image $\mathbf{x}_{t-1}$. Repeating this process over $T$ steps produces a noise-free image $\mathbf{x}_{0}$.
Operating directly on the image $\mathbf{x}_{t}$ in pixel-space is computationally expensive. As an alternative, latent diffusion models have been proposed to curtail this high cost, in which the denoising procedure is performed in latent space, whose dimensionality is typically much lower than pixel space. Stable Diffusion Rombach et al. (2022) is one example of a latent diffusion model which achieves state-of-the-art performance on various image synthesis tasks. It leverages a powerful autoencoder to project to and from latent space, where the standard denoising procedure is performed. Images in latent space are typically denoted by $\mathbf{z}_{t}$, and the encoder and decoder are denoted by $\mathcal{E}$ and $\mathcal{D}$, respectively, $\mathbf{z}_{t}=\mathcal{E}(\mathbf{x}_{t})$ and $\mathbf{x}_{t}=\mathcal{D}(\mathbf{z}_{t})$.
During training, samples from the true data distribution $q(\mathbf{x}_{0})$ are corrupted via the forward process. By training a diffusion model to learn a reverse process in which it iteratively reconstructs these noisy samples into noise-free samples, it is possible to generate images from pure noise at inference time. This corresponds to sampling from an approximation $p_{\theta}(\mathbf{x}_{0})$ to the data distribution, $q(\mathbf{x}_{0})$. This generation process can be guided by introducing an additional input vector $\mathbf{y}$, which is often a text prompt. In this case, the model produces samples from an approximation $p_{\theta}(\mathbf{x}_{0}|\mathbf{y})$ to the conditional distribution $q(\mathbf{x}_{0}|\mathbf{y})$.
In denoising diffusion probabilistic models (DDPM) Ho et al. (2020), the forward process is characterized by the Markov chain $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})\sim\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$, for some noise schedule $\beta_{t}$. In this case, $q(\mathbf{x}_{t}|\mathbf{x}_{0})\sim\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ and $\alpha_{t}=1-\beta_{t}$. The reverse process is typically modeled by a learned Markov chain $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\sim\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\sigma_{t}^{2}\mathbf{I})$, where $\sigma_{t}$ is an untrained time dependent constant, usually with $\tilde{\beta}_{t}\leq\sigma_{t}^{2}\leq\beta_{t}$ and $\tilde{\beta}_{t}=\beta_{t}(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})$, or with $\sigma_{t}$ simply chosen equal to $\sqrt{\beta_{t}}$.
@@
The right hand side is the usual evidence lower bound (ELBO), which is minimized instead. Ho et al. Ho et al. (2020) show that minimizing the ELBO is equivalent to minimizing
@@
for some positive function $\lambda(t)$, where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $\mathbf{x}_{t}(\mathbf{x}_{0},\epsilon)=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, and